OpenStudy (futureboy1):

Simplify the expression. 3 • (7 + 2)3

pooja195 (pooja195):

Use PEMDAS 

OpenStudy (futureboy1):

I feel the answer is 19,683.....am I right?

pooja195 (pooja195):

Not even close lol Parentheses come first lets add 7+2 7+2=?

OpenStudy (futureboy1):

9

pooja195 (pooja195):

Good now all we have to do is multiply \[\huge ~\rm~\bf (3)(9)(3)\]

pooja195 (pooja195):

\[\huge~\rm~3 \times 9 \times 3=?\]

OpenStudy (futureboy1):

But 9*3=27 and then it's to the 3rd power so isn't it 27*27*27?

pooja195 (pooja195):

O_o to the ^3? Wait so your question looked like this \[\huge~\rm~3 • (7 + 2)3^3? \]

OpenStudy (futureboy1):